$SPAD/src/input array1.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool array1.output
)set message test on
)set message auto off
)clear all
--S 1 of 9
oneDimensionalArray [i**2 for i in 1..10]
--R
--R
--R   (1)  [1,4,9,16,25,36,49,64,81,100]
--R                                   Type: OneDimensionalArray(PositiveInteger)
--E 1
--S 2 of 9
a : ARRAY1 INT := new(10,0)
--R
--R
--R   (2)  [0,0,0,0,0,0,0,0,0,0]
--R                                           Type: OneDimensionalArray(Integer)
--E 2
--S 3 of 9
for i in 1..10 repeat a.i := i; a
--R
--R
--R   (3)  [1,2,3,4,5,6,7,8,9,10]
--R                                           Type: OneDimensionalArray(Integer)
--E 3
--S 4 of 9
map!(i +-> i ** 2,a); a
--R
--R
--R   (4)  [1,4,9,16,25,36,49,64,81,100]
--R                                           Type: OneDimensionalArray(Integer)
--E 4
--S 5 of 9
reverse! a

--R
--R
--R   (5)  [100,81,64,49,36,25,16,9,4,1]
--R                                           Type: OneDimensionalArray(Integer)
--E 5
--S 6 of 9
swap!(a,4,5); a
--R
--R
--R   (6)  [100,81,64,36,49,25,16,9,4,1]
--R                                           Type: OneDimensionalArray(Integer)
--E 6
--S 7 of 9
sort! a
--R
--R
--R   (7)  [1,4,9,16,25,36,49,64,81,100]
--R                                           Type: OneDimensionalArray(Integer)
--E 7
--S 8 of 9
b := a(6..10)
--R
--R
--R   (8)  [36,49,64,81,100]
--R                                           Type: OneDimensionalArray(Integer)
--E 8
--S 9 of 9
copyInto!(a,b,1)
--R
--R
--R   (9)  [36,49,64,81,100,36,49,64,81,100]
--R                                           Type: OneDimensionalArray(Integer)
--E 9
)spool
)lisp (bye)

———-

References^{[1] nothing}

5